OpenStudy (lena772):

(a) find an equation of the tangent line to the graph of at the given point, (b) use a graphing utility to graph the function and its tangent line at the point, and (c) use the derivative feature of the graphing utility to confirm your results.

OpenStudy (lena772):

function: y= cos 3x point: (pi/4, -sqrt2/2)

ganeshie8 (ganeshie8):

where wer u stuck ?

ganeshie8 (ganeshie8):

i saw u doing a similar problem the other day ? :)

OpenStudy (anonymous):

take the derivative

OpenStudy (lena772):

I just get confused when trig is added in

OpenStudy (anonymous):

do you know what the derivative is

OpenStudy (lena772):

dy/dx

OpenStudy (anonymous):

-3sin3x

OpenStudy (anonymous):

right

OpenStudy (lena772):

yes

ganeshie8 (ganeshie8):

so, slope = m = dy/dx = -3sin3x

ganeshie8 (ganeshie8):

since u want the slope at point: (pi/4, -sqrt2/2) evaluate slope at x = pi/4

ganeshie8 (ganeshie8):

dy/dx at x = pi/4 :- -3sin(3*pi/4) = ?

OpenStudy (lena772):

-0.123

OpenStudy (lena772):

@ganeshie8

ganeshie8 (ganeshie8):

that doesnt look right you should get :- dy/dx at x = pi/4 :- -3sin(3*pi/4) = \(\large \frac{-3}{\sqrt{2}}\)

ganeshie8 (ganeshie8):

now that u knw slope, m = \(\large \frac{-3}{\sqrt{2}}\) point: (pi/4, -sqrt2/2) wats the equation of tangent in point slope form ?

OpenStudy (lena772):

y-(-sqrt2/3)=(-3/sqrt2)(x-(pi/4)

OpenStudy (lena772):

?

OpenStudy (lena772):

@hartnn

ganeshie8 (ganeshie8):

thats right !

OpenStudy (lena772):

So that's for a? What would the graph look like for b? Do I just put that equation into Geogebra?

OpenStudy (lena772):

@ganeshie8

ganeshie8 (ganeshie8):

yup graph all 3 below :- 1) function: y= cos 3x 2) point: (pi/4, -sqrt2/2) 3) tangent : y-(-sqrt2/\(\color{red}{2}\))=(-3/sqrt2)(x-(pi/4)

ganeshie8 (ganeshie8):

u should get exactly like the one attached